# Hierarchical Text Classification  (HTC) vs. eXtreme Multilabel Classification (XML): Two Sides of the Same Medal

### v2

@@ Abstract.
Assigning a subset of labels from a fixed pool of labels to a given input text is a text classification problem with many real-world applications, such as in recommender systems. Two separate research streams address this issue. Hierarchical Text Classification (HTC) focuses on datasets with smaller label pools of hundreds of entries, accompanied by a semantic label hierarchy. In contrast, eXtreme Multi-Label Text Classification (XML) considers very large label pools with up to millions of entries, in which the labels are not arranged in any particular manner. However, in XML, a common approach is to construct an artificial hierarchy without any semantic information before or during the training process. Here, we investigate how state-of-the-art models from one domain perform when trained and tested on datasets from the other domain. The HBGL and HGLCR models from the HTC domain are trained and tested on the datasets Wiki10-31K, AmazonCat-13K, and Amazon-670K from the XML domain. On the other side, the XML models CascadeXML and XR-Transformer are trained and tested on the datasets Web of Science, The New York Times Annotated Corpus, and RCV1-V2 from the HTC domain. HTC models, on the other hand, are not equipped to handle the size of XML datasets and achieve poor transfer results. The code and numerous files that are needed to reproduce our results can be obtained from https://github.com/FloHauss/XMC_HTC.
NLP, text classification, hierarchical text classification, extreme multi-labeling
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Document analysisccs: Computing methodologies Information extraction
1. Introduction
Text classification is an essential task in text mining and natural language processing (NLP), which involves assigning pre-defined labels to input text, with applications ranging from recommender systems and information retrieval to tasks such as sentiment analysis, e-commerce, advertising and news classification (Gasparetto et al., 2022; Chang et al., 2021; Jain et al., 2019; Sun and Lim, 2001a). Since the introduction of text classification decades ago, many approaches have been developed, with a recent trend focusing on deep learning models (Li et al., 2020). Within the landscape of multi-label text classification, two main categories of approaches emerged: Hierarchical Text Classification (HTC) and eXtreme Multi-Label Text Classification (XML). Both approaches can be seen as distinct and relevant methodologies addressing specific issues and challenges in text classification.
@@ 1. Introduction
HTC is a text classification problem tailored to exploit the inherent information of a hierarchically structured labelset (Zhou et al., 2020a). Furthermore, HTC utilizes a top-to-bottom approach for path construction and hierarchical relationships to improve classification performance (Sun and Lim, 2001b). A common approach to taxonomic hierarchy in HTC is to model a graph or tree, where each node represents a label to be classified (Wang et al., 2022a). In the field of HTC, there have been many different approaches that incorporate hierarchical information. For instance, independently processing text and structure of label hierarchy and then merging the two elements in a joint representation (Chen et al., 2021a). By doing so, it is aimed to reach a combined representation of text and hierarchy that is beneficial for classification (Zhou et al., 2020a). Another usage of hierarchy for text classification within HTC can be narrowed down to local and global approaches  (Zhou et al., 2020b), where local approaches aim to create classifiers at each level of the hierarchy by using strategies such as CNNs. On the other hand, global strategies approach HTC as a form of multi-label classification by integrating hierarchical information directly into models to rely on the inherent structure and relationships within the hierarchy to improve classification accuracy (Sun and Lim, 2001a).
@@
Unlike HTC, XML is designed to handle large datasets and to work with a large number of labels. XML is an advanced text classification domain that focuses on assigning a relevant subset of instances from a large labelset, where the number of labels poses serious computational and statistical challenges, making it difficult to create an XML method that is efficient to train and accurate (Zhang et al., 2024; You et al., 2019). XML methods can be broadly categorized into three areas: sparse linear models, embedding-based methods, and transformer-based models (Zhang et al., 2024). Initially, sparse linear models laid the foundation. They use linear algorithms and efficient partitioning to work with extensive label spaces by implementing hierarchical trees and multi-label classification techniques such as one-versus-all (Wydmuch et al., 2018). Subsequently, embedding-based techniques were introduced to work with neural networks and aimed to learn semantic representations, often combined with sampling for scalability (You et al., 2019). The latest development, transformer-based models, exploits the capabilities of pre-trained transformers to capture complex semantic relationships within the data (Zhang et al., 2024).
Motivation
@@
So far, it is unclear how the XML and HTC techniques would perform in unconventional settings, namely when XML methods encounter hierarchical data and HTC methods work with vast amounts of unstructured labels. Driven by a cross-domain evaluation, this study aims to identify potential limitations and opportunities for improvement within HTC and XML approaches and to assess ways in which the methodologies can be extended to accommodate more diverse datasets. The investigation is expected to provide key information regarding the performance and usability of HTC and XML approaches and potential new applications in the field of multi-label text classification.
Problem Statement
@@
Until now, both HTC and XML have co-existed as two distinct domains of research in the world of multi-label text classification. Both approaches aim to train a classifier on a given dataset containing texts and their respective class labels to predict relevant labels for an input. However, what distinguishes these two approaches is, firstly, that HTC models are additionally provided with a fixed label taxonomy. XML methods create such a hierarchy through clustering at runtime. Secondly, the number of labels in the HTC datasets is much smaller than the number of labels in the XML datasets. A few hundred in the case of HTC datasets, but up to millions when it comes to XML. For that reason, there has been no research or knowledge about a model performance, when applied to datasets from the other domain. To gain a deeper understanding of the individual capabilities of state-of-the-art models from both sides, this separation has to be dissolved by making the datasets from one side accessible to models from the other, without structural information loss or unfeasible spatial or computational costs. A further distinction is that the evaluation measures differ between the HTC domain and the HTC domain, HTC commonly employs standard classification metrics such as F1-micro and F1-macro, whereas XML uses ranking-based metrics such as P@k.
@@
We have developed a novel approach that allows cross-domain comparison of HTC and XML models with datasets traditionally used for text classification tasks from the other side. This signifies that we have transferred HTC models to be applicable to XML datasets and vice versa. Furthermore, the R-Precision metric has been applied to all models to ensure a fair comparison. Our results show that XML models, specifically CascadeXML, achieve comparable or better results than HTC models when tested on datasets from both domains for the performance measures P@k and R-Precision. This allows us to argue that XML models are more versatile. The results of our study suggest that models need to be tested more frequently in cross-domain tasks and that at least XML models need to be considered as baselines in HTC research.
Organisation
In Section 2 we summarize related work, subdivided into HTC and XML. In Section 3 we describe how we bring the two worlds together. Here we focus on the modification of benchmark datasets from one world in order to use them in the other. The experimental apparatus is described in Section 4, where preprocessing steps, the general procedure, hyperparameter optimization, and performance measures are explained. An overview of the obtained results is given in Section 5, which are then discussed in Section 6.
2. Related Work
@@
For both, the single and multi-label text classification tasks, fine-tuning pre-trained transformer models yield state-of-the-art performance (Li et al., 2020). They outperform generative large language models such as GPT-4 (OpenAI, 2023) and Llama-2 (Touvron et al., 2023) in accuracy while offering advantages of smaller model size and less training time (Sun et al., 2023; Yu et al., 2023). They also improve on recent graph-based methods (Galke et al., 2022). Therefore, we focus on models based on pre-trained transformers as well as common baseline methods. Below, we discuss the state-of-the-art in both domains separately, first in HTC, followed by XML. For further reading, we refer to surveys like Zangari et al. (2023).
2.1. Hierarchical Text Classification (HTC)
@@
Given a hierarchy of class labels organized as a tree (in short, a label tree) and an input text, the task of HTC is to produce a set of labels from the label tree that accurately represents the input text. Approaches in this field are often distinguished as local and global (Zhou et al., 2020b), where local approaches focus on building appropriate classifiers for each node or each level in the label tree, whilst global approaches employ a single classifier for the whole hierarchy instead. The models we selected generally follow the global approach.
HGCLR (Wang et al., 2022a) is a recent model that integrates the label hierarchy directly into the text encoder using a contrastive learning framework. It employs BERT (Devlin et al., 2019) as the text encoder, which processes the input token sequence and produces hidden representations for each token, with the [CLS] token summarizing the entire sequence for classification tasks. As a graph encoder, HGCLR employs a modified Graphormer (Ying et al., 2021) to capture hierarchical structures. The modified Graphormer integrates spatial encoding, edge encoding, and centrality encoding into its architecture. The self-attention mechanism of this modified Graphormer is expressed as follows:
${\bm{A}}_{{\bm{G}}_{ij}}=\frac{({\bm{f}}_{i}{\bm{W}}_{\bm{Q}}^{\bm{G}})({\bm{f}}_{j}{\bm{W}}_{\bm{K}}^{\bm{G}})^{\top}}{\sqrt{{\bm{d}}_{h}}}+c_{ij}+{\bm{b}}_{\phi(y_{i},y_{j})}$
[…]
HBGL (Jiang et al., 2022) builds upon the ideas of HGCLR. The key difference is that HBGL removes the auxiliary graph encoder component and uses BERT (Devlin et al., 2019) for both, text encoding and graph feature encoding. The model consists of two separate training stages using the same BERT model. In the first one, the global hierarchy-aware label embedding module, the label name is tokenized and used to inject the global hierarchy of the hierarchical label tree to create initial label embeddings. Within this step, the BERT encoder is frozen. Learning is performed by using a modified version of the BERT’s masked language modeling (MLM) objective and by using a custom attention mask to represent the hierarchy. The masked label prediction task is treated as multi-label classification, where a masked leaf label must predict itself and its masked sibling leaf labels. This approach prevents BERT from confusing masked leaf labels under the same parent label. In the second stage, which is the local hierarchy-aware text encoder, the BERT encoder receives three inputs: The appropriate text tokens, the local hierarchy, and the masked local hierarchy. The local hierarchy consists of one token for each hierarchy level, representing an aggregation of the relevant label embeddings within the respective hierarchical level. Likewise, the masked local hierarchy consists of additional masked tokens, one for each of the hierarchy tokens. The model learns to predict this local hierarchy in an autoregressive fashion. Learning is done in the manner of s2s-ft (Bao et al., 2021), again with a custom attention mask to prevent higher levels in the hierarchy from attending lower levels. The predicted local hierarchy can then be used for inference.
Other models that leverage hierarchical information in the context of text classification and work similarly to HGCLR, such as HPT (Wang et al., 2022b), use different graph encoders, such as GAT (Velickovic et al., 2018), and different learning approaches. For example, they focus more on MLM rather than contrastive learning. Certain methods also do not rely on pre-trained transformer models like BERT and instead utilize transformer mechanisms such as attention, as seen in HiAGM (Zhou et al., 2020b). However, research has demonstrated that leveraging pre-trained transformers generally enhances performance. This stance is supported by findings presented by Galke et al. (2022), and further asserted by earlier studies such as HTCInfoMax (Deng et al., 2021) and HiMatch (Chen et al., 2021b), which have explored various methodologies for integrating textual and hierarchical data. Overall, pre-trained transformer-based models seem to achieve state-of-the-art results in the field of HTC. From the mentioned models we select HBGL and HGCLR for our evaluation as both of them use the same datasets and metrics, and show the best performance.
2.2. eXtreme Multilabel Classification (XML)
@@
XML, like HTC, aims to assign a subset of labels to a given text and leverages a hierarchy to improve accuracy and reduce training complexity. However, there are two main differences: Firstly, the hierarchies are not given as additional input but are created by the model itself based on vector representations of the labelset. Secondly, the XML label spaces are much larger than their HTC counterparts.
@@
There are two avenues in which XML techniques can be grouped based on input: Firstly, there are conventional machine learning methods that employ sparse features such as bag-of-words (BoW) features from text. Secondly, there are deep learning methods utilizing the raw textual data. XML algorithms in the first category can be divided into the following three subclasses: one-vs-all (Babbar and Schölkopf, 2017; Yen et al., 2017), label-tree-based (Prabhu et al., 2018), and embedding-based (Bhatia et al., 2015) methods. Recently, deep learning methods from the second category that operate on the full sequence of text inputs (rather than a bag of words) have become popular. They fine-tune pre-trained transformer models like BERT, RoBERTa (Liu et al., 2019), or XLNet (Yang et al., 2019) and have shown promising results in both prediction accuracy and model size. In the following, we will discuss the architectures and strengths and weaknesses of some recent approaches.
AttentionXML (You et al., 2019) as a label-tree-based deep learning model combines two such techniques to take advantage of both approaches. It does not yet use the transformer architecture, instead a BiLSTM (Schuster and Paliwal, 1997) is used. In addition to that, AttentionXML employs two features: The first is a multi-label attention layer, to incorporate different contextual information from the raw input for each label, alongside the semantic information provided by using raw text instead of bag-of-words features. The second one is the creation and usage of a shallow and wide probabilistic label tree for its training, to be able to handle the enormous label spaces present in XML tasks and thus improve the scalability of datasets with a large pool of labels. Through its usage of said tree and the multi-label attention, AttentionXML also shows good performance in predicting rare labels.
To improve efficiency and accuracy LightXML (Jiang et al., 2021) employs end-to-end training and dynamic negative label sampling via generative cooperative networks, allowing the label ranking part to learn progressively from easy to hard instances, preventing overfitting and improving overall model performance. Unlike static negative sampling methods such as AttentionXML, the dynamic negative sampling approach ensures that negative labels are re-sampled for each training instance. LightXML also makes use of pre-trained transformers to obtain text embeddings. The base models of BERT, RoBERTa, and XLNet are used to generate multi-layer text representations by concatenating the hidden state of the [CLS] token of the last five layers of the encoder, which speeds up convergence. Although LightXML outperforms AttentionXML in terms of model size and training time, it scales poorly compared to newer models described below, especially on large datasets.
@@
To address the issue of scalability and the long training times required for the largest datasets, XR-Transformer (Zhang et al., 2024) introduces a recursive architecture. Instead of selecting a fixed label resolution for training, the pre-trained transformer model is gradually refined by increasing label resolutions. This creates a series of simple to difficult training tasks and reduces training costs. To accomplish this, a Hierarchical Label Tree (HLT) is formed by repeatedly applying a $k$-means clustering algorithm on the label features. The feature of a label $l\in L$ is computed by averaging the TF-IDF (Term Frequency-Inverse Document Frequency) representations of the documents that carry a label $l$. Subsequently, the model is trained on each level of the tree in a top-down manner. This multi-resolution learning approach is also studied in other machine learning contexts, such as computer vision (Zhou and Liang, 2023). At each level, a three-step algorithm is executed. The first step, shortlisting, involves selecting the top-$k$ clusters from the parent level. These clusters are then used in the second step, fine-tuning, where the model is fine-tuned on the label resolution of the current level. Finally, in the third step, ranking, the top-$k$ clusters from the current level are ranked. XR-Transformer has been demonstrated to reduce the training times of LightXML by an average factor of 17 while maintaining similar accuracy results.
@@
CascadeXML (Kharbanda et al., 2022) was introduced to address the complexity of XR-Transformer and the scalability of LightXML. It is an end-to-end trainable XML algorithm that outperforms the aforementioned approaches on the P@k metric. The model uses the same ensemble settings as LightXML and XR-Transformer. It utilizes an ensemble of BERT, RoBERTa, and XLNet for smaller datasets and an ensemble of three BERT models with different seeds for random initialization for larger datasets. CascadeXML takes advantage of the multi-layer architecture of transformers and does not treat them as black-box encoders. Instead, CascadeXML trains a single transformer across multiple resolutions, allowing it to use attention maps specific to the tree resolution. At its core, the method clusters all the given labels and creates an HLT – similar to the XR-Transformer. A shortlisting process is performed to select the highest-scoring metalabels, which are referred to as level-one metalabels. This process is continued until the final level is reached, which, in contrast to XR-Transformer, carries out classification at full label resolution.
Inspired by InceptionNet, InceptionXML (Kharbanda et al., 2023) proposes a lightweight architecture for extreme short text classification. It uses an embedding enhancement module to make representations independent of word order. The architecture employs three parallel branches of 1D convolution layers with different filter sizes, followed by an extreme linear classifier. The model integrates skip connections, non-linear activation, and one-vs-all classifiers for label prediction. Training involves end-to-end optimization using binary cross-entropy loss. Overall, InceptionXML achieves state-of-the-art results in extreme multi-label classification for short text queries.
2.3. Transferability
Although HTC and XML both work with multiple labels, HTC methods require input labels to be structured in a hierarchical tree. Therefore, HTC and XML use completely different benchmark datasets. On the HTC side, widely used benchmark datasets are Web of Science (WoS) (Kowsari et al., 2017), The New York Times Annotated Corpus (NYT) (Shimura et al., 2018), and RCV1-v2 (Lewis et al., 2004). On the other hand, Wiki10-31K (Zubiaga, 2012), AmazonCat-13K (McAuley and Leskovec, 2013), and Amazon-670K (McAuley and Leskovec, 2013) are commonly used in the XML literature. Note that these datasets are different in size and evaluation strategy for HTC and XML models. HTC models use datasets with smaller labelsets and fewer data points than XML datasets and evaluate performance by calculating Micro-F1 and Macro-F1 scores. XML models are evaluated with ranking-based metrics, such as P@k and PSP@k (Jain et al., 2016) metrics. This also poses issues related to the comparability of the two worlds. Micro-F1 and Macro-F1 scores are calculated based on hard yes or no classification decisions. P@k and PSP@k are ranking-based metrics that have a somewhat arbitrary cut-off point after the first $k$ predictions, which hinders computing a reasonable F1 score. In order to conduct fair experiments, we implemented the P@k metric for the HTC models and applied the R-Precision metric (Manning et al., 2009) for models from both domains to ensure that the results can be adequately compared. This is described in detail in Section 4.5.
3. Dataset Transfer XML $\leftrightarrow$ HTC

### v3

# Your Extreme Multi-label Classifier is Secretly a Hierarchical Text Classifier for Free

@@ Abstract
Assigning a set of labels to a given text is a classification problem with many real-world applications, such as recommender systems. Two separate research streams address this issue. Hierarchical Text Classification (HTC) focuses on datasets with label pools of hundreds of entries, accompanied by a semantic label hierarchy. In contrast, eXtreme Multi-Label Text Classification (XML) considers very large sets of labels with up to millions of entries but without an explicit hierarchy. In XML methods, it is common to construct an artificial hierarchy in order to deal with the large label space before or during the training process. Here, we investigate how state-of-the-art HTC models perform when trained and tested on XML datasets and vice versa using three benchmark datasets from each of the two streams. Our results demonstrate that XML models, with their internally constructed hierarchy, are very effective HTC models. HTC models, on the other hand, are not equipped to handle the sheer label set size of XML datasets and achieve poor transfer results. We further argue that for a fair comparison in HTC and XML, more than one metric like F1 should be used but complemented with P@k and R-Precision. Source code is available on EasyChair.
1 Introduction
Text classification is an essential task in text mining and natural language processing (NLP), which involves assigning pre-defined labels to input text, with applications in recommender systems, information retrieval, sentiment analysis, and more [9, 4, 11, 27]. Since the introduction of text classification decades ago, many approaches have been developed, with a recent trend focusing on deep learning models [19]. When considering more than one possible output per example, a.k.a. multi-label classification, the literature is split into two streams. On the one hand, there is hierarchical text classification (HTC), which aims to make use of a hierarchical structure among the classes [32, inter alia]. On the other hand, there is extreme multi-label text classification (XML), which aims to develop techniques to handle an extensive set of possible classes [38, inter alia]. From a conceptual point of view, both families of approaches aim to tackle the same problem: multi-label text classification. However, the literature is split in the sense that researchers employ different datasets and different evaluation metrics and do not consider the methods from the other side for comparison when developing new methods. Interestingly, XML methods often use an artificial hierarchical structure to deal with the large label space. We question whether this harsh distinction of problems and methods can be alleviated and experiment with methods from XML applied to datasets used in HTC and vice-versa.
@@ 1 Introduction
In HTC, the text classification models exploit the inherent information of a hierarchically structured set of classes [44]. Furthermore, HTC utilizes a top-to-bottom approach for path construction and hierarchical relationships to improve classification performance [27]. A common approach to the taxonomic hierarchy in HTC is to model a graph or tree, where each node represents a label to be classified [32]. In the field of HTC, there have been many different approaches that incorporate hierarchical information. For instance, independently processing text and structure of label hierarchy and then merging the two elements into a joint representation [5]. By doing so, it is aimed to reach a combined representation of text and hierarchy that is beneficial for classification [44]. Another usage of hierarchy for text classification within HTC can be narrowed down to local and global approaches [44], where local approaches aim to create classifiers at each level of the hierarchy by using strategies such as CNNs. On the other hand, global strategies approach HTC as a form of multi-label classification by integrating hierarchical information directly into models to rely on the inherent structure and relationships within the hierarchy to improve classification accuracy [27].
@@
In XML, the main objective is to be able to handle large datasets with a large number of labels. XML is an advanced text classification domain that focuses on assigning a relevant subset of labels to an instance, where the number of labels poses serious computational and statistical challenges, making it difficult to create an XML method that is efficient to train and accurate [42, 38]. XML methods can be broadly categorized into sparse linear models, embedding-based methods, and transformer-based models [42]. Initially, sparse linear models laid the foundation. They use linear algorithms and efficient partitioning to work with extensive label spaces by implementing hierarchical trees and multi-label classification techniques such as one-versus-all [34]. Subsequently, embedding-based techniques were introduced to work with neural networks and aimed to learn semantic representations, often combined with sampling for scalability [38]. Recent developments make use of pre-trained transformer language models to capture complex semantic relationships within the data [42].
@@
So far, it is unclear how the XML and HTC techniques would perform in unconventional settings, namely when XML methods encounter hierarchical data and HTC methods work with vast amounts of labels. Driven by a cross-domain evaluation, this study aims to identify opportunities for improvement within HTC and XML approaches and potential limitations. We expect to gain key insights on the performance and usability of HTC and XML approaches and potential new applications in the field of multi-label text classification.
@@
Until now, HTC and XML have existed as two distinct streams of research in multi-label text classification. Both streams aim to train a classifier on a given dataset containing texts and their respective class labels to predict relevant labels for an input. What distinguishes these two approaches is, firstly, that HTC models are additionally provided with a fixed label taxonomy. XML methods create such a hierarchy through clustering at runtime. Secondly, the number of labels in the HTC datasets is much smaller than in the XML datasets. A few hundred in the case of HTC, but up to millions when it comes to XML. For that reason, there has been no research or knowledge about model performance when applied to datasets from the other domain. To gain a deeper understanding of the individual capabilities of state-of-the-art models from both sides, this separation has to be dissolved by making the datasets from one side accessible to models from the other side without structural information loss or unfeasible computational costs. A further distinction is that the evaluation measures differ between the HTC domain and the XML domain. HTC commonly employs Micro-/Macro-F1, while XML uses ranking-based metrics such as P@k.
@@
So far, it was unclear whether the XML methods would be applicable to HTC datasets. We contribute a procedure on how to overcome the barriers of different dataset types and metrics, allowing future researchers and practitioners to employ XML methods for HTC datasets and vice versa. To facilitate such a comparison, we have harmonized the metrics and suggest using R-precision. We show how the models can be applied to the datasets of the respective other research stream. Our results show that XML models, specifically CascadeXML, achieve comparable or better results than HTC models when tested on datasets from both domains for the performance measures P@k and R-Precision. This allows us to argue that XML models are more versatile. Our results suggest that XML and HTC are closer to each other than one would expect on the basis of a clearly divided literature. Our main finding is that XML models – despite not making use of the manually curated hierarchies of HTC datasets – are competitive to state-of-the-art HTC methods.
2 Related Work
@@
Generally, the state of the art in single and multi-label text classification is fine-tuning a pre-trained encoder-only language model [19, 8] such as BERT [7]. These models even outperform generative large language models such as GPT-4 [23] and Llama-2 [30] in accuracy while offering advantages of smaller model size and less training time [28, 39]. They also improve on recently proposed graph-based methods [8]. Therefore, we focus on models based on pre-trained transformers as well as common baseline methods. We first discuss the state-of-the-art in HTC followed by XML. For further reading, we refer to surveys like Zangari et al. [2023].
Hierarchical Text Classification (HTC)
@@
Given a hierarchy of class labels organized as a tree (in short, a label tree) and an input text, the task of HTC is to produce a set of labels from the label tree that accurately represents the input text. Approaches in HTC are often distinguished as local and global [44], where local approaches focus on building appropriate classifiers for each node or each level in the label tree, whilst global approaches employ a single classifier for the whole hierarchy. The state-of-the-art models HGCLR and HBGL follow the global approach.
HGCLR [32] is a model that effectively integrates the label hierarchy with the semantics from the text and labels using a contrastive learning framework. It employs BERT [7] as the text encoder and a modified Graphormer [37] to capture hierarchical structures. The key distinction in HGCLR’s modified Graphormer is the combination of a label embedding and a name embedding. The former is an embedding of the vertex in the hierarchy, and the latter is a textual embedding of the label string.
HBGL [13] removes the auxiliary graph encoder and uses BERT [7] for both text and graph encoding. The model consists of two separate training stages using the same BERT model. In the first stage, the approach learns label embeddings based on random walks through the hierarchy, while keeping the BERT model frozen. Learning is performed using a modified version of BERT’s masked language modeling objective and a custom attention mask to represent the hierarchy. In the second stage, and based on these initial label vectors, the BERT model is trained to sequentially generate class labels based on the text input with an objective similar to sequence-to-sequence fine-tuning [2].
Other models that leverage hierarchical information that work similarly to HGCLR are HPT [33], which employs different graph encoders such as graph attention [31], and HALB [43], which replaces the classification loss with an asymmetric loss and adds another loss for separating samples with similar representations but different labels. HILL [45] applies a hierarchy-aware contrastive loss, too, using BERT as a text encoder and a separate graph encoder. Earlier methods did not rely on transformer models, such as HiAGM [44]. However, leveraging pre-trained transformers has generally shown to enhance performance [8, 6, 5] and resembles the state of the art in HTC with HBGL [13].
eXtreme Multi-label Classification (XML)
@@
XML, like HTC, aims to assign a subset of labels to a given text and leverages a hierarchy to improve accuracy and reduce training complexity. However, there are two main differences: First, the label spaces usually are much larger than in HTC. Second, the hierarchies do not come from the data as in HTC but are artificially constructed in XML.
@@
There are two main aspects along which XML techniques can be categorized: First, there are conventional machine learning methods that employ sparse features such as bag-of-words (BoW) features from text. Second, there are deep learning methods that operate on the raw text data. XML algorithms in the first category can be divided into the following three subcategories: one-vs-all [1, 36], label-tree-based [24], and embedding-based [3] methods. Recently, deep learning methods from the second category that operate on the full sequence of text inputs (rather than a bag of words) have become increasingly popular [16, 42]. They fine-tune pre-trained transformer models like BERT, RoBERTa [20], or XLNet [35] and have shown promising results in both prediction accuracy and model size. Strong performers in XML classification, among others, are AttentionXML [38] using BiLSTM [25] and InceptionXML [15] for extreme short text classification. LightXML [14] employs an ensemble of BERT, RoBERTa, and XLNet together with dynamic negative label sampling via generative cooperative networks.
@@
We discuss in detail the two strongest models for XML, which we consider for our study. The XR-Transformer [42] is a recursive architecture where the transformer layers are connected with a gradually refined increase in label resolutions. To accomplish this, a hierarchical label tree is formed by repeatedly applying a $k$-means clustering algorithm on the label features. The feature of a label $l\in L$ is computed by averaging the TF-IDF representations of the documents that carry a label $l$. Subsequently, the model is trained on each level of the tree in a top-down manner. At each level, a three-step algorithm is executed. The first step involves selecting the top-$k$ clusters from the parent level. These clusters are then used in the second step, where the model is fine-tuned on the label resolution of the current level. Finally, the top-$k$ clusters from the current level are ranked. XR-Transformer uses the same ensemble as LightXML but is 17x faster with similar accuracy.
@@
CascadeXML [16] addresses the complexity of XR-Transformer and outperforms the aforementioned approaches on the P@k metric. It uses an ensemble similar to XR-Transformer and LightXML. CascadeXML takes advantage of the multi-layer architecture of transformers and does not treat them as black-box. Instead, CascadeXML trains a single transformer across multiple resolutions, allowing it to use attention maps specific to the current tree resolution. The method clusters all given labels and creates an HLT as in XR-Transformer. A shortlisting process is performed to select the highest-scoring meta-labels, which are referred to as level-one meta-labels. This process is continued until the final level is reached, which, in contrast to XR-Transformer, carries out classification at full label resolution.
3 Transfer XML $\leftrightarrow$ HTC
Due to the different nature of tasks tackled, HTC and XML studies use different benchmark datasets. On the HTC side, widely used benchmark datasets are Web of Science [17], The New York Times Annotated Corpus [26], and RCV1-v2 [18]. On the other hand, Wiki10-31K [46], AmazonCat-13K [22], and Amazon-670K [22] are commonly used in the XML literature. Note that these datasets differ in size, number of classes, and evaluation strategy for HTC and XML models. HTC models use datasets with smaller label sets and fewer data points than XML datasets and evaluate performance by calculating the classification metrics Micro-F1 and Macro-F1 scores. In contrast, XML models are commonly evaluated with ranking-based metrics, such as P@k and PSP@k [12] metrics. This poses challenges regarding the comparability of the two streams of research.
In more detail, Micro-F1 and Macro-F1 scores are calculated based on hard yes or no classification decisions. P@k and PSP@k are ranking-based metrics that have an arbitrary cut-off point (as specified by $k$) after the highest-ranking $k$ predictions. This hinders computing an F1 score over the full range of classes. For a fair comparison, we decided to employ the P@k metric for the HTC models and additionally consider the R-Precision metric [21] for models from both domains to ensure that the results can be fairly compared. We describe this in more detail in Section 3.2.
3.1 Bidirectional Transfer of Datasets